I need help in these questions: please help !

1. A balloon containing 157 moles of ammonia gas has an initial volume of 60.0 mL at a temperature of 80.0 C. What is its final temperature when the volume is reduced to 50 mL?

2. An atmospheric sample contains nitrogen at 554 torr, carbon dioxide at 170 torr and oxygen. Assuming STP conditions and a total of 10.0 moles of the gas mixture, what is the partial pressure of oxygen in torr?

2b. How many moles of oxygen are present in this sample?

3. 50.15 grams if Xenon gas (Xe) occupy a volume of 2.80 L at a pressure of 200 torr. What is the temperature of the gas in Kelvin?
Answer #1

1)                    Initial condition                                                         Final condition

Volume = V1= 60.0 mL                                             volume = V2= 50.0 mL

Temperature = T1= 80.0C=80.0+273=353 K           Temperature = T2 =

according to charles law,

At constant pressure , given mass of a gas,volume is directly proportional to the absolute temperature.

V is proprotional to the Temperature.

V1/T1 = V2/T2

60.0/353 = 50.0/T2

T2= 294.17K

Final temperature = 294.17K

3)

mass of Xe= 50.15 grams

molar mass of Xe = 131.29 gram/mole

number of moles of Xe = mass/molar mass = 50.15/131.29 = 0.3819 mole

number of moles = n= 0.3819 moles

volume = 2.80L

Pressure = 200 torr= 0.263 atm              [ 1 torr = 0.00131579 atm ]

Gas constant = R = 0.0821 L-atm/mol-K

PV=nRT

T = PV/nR

T = 0.263 x 2.80/0.3819 x0.0821

T= 23.49 K